[figure: 9.7 cm]

PSEUDOGAP AND PAIRING DUE TO KINETIC ENERGY

SAVING IN DOPED ANTIFERROMAGNETS

P. Wróbel^{b}, R. Eder^{a}, and R.Micnas^{c} aInstitute for Low Temperature and Structure Research P. 0. Box 1410, 50-950 Wrocław 2, Poland bInstitut für Theoretische Physik, Universität Würzburg Am Hubland, 97074 Würzburg, Germany c Adam Mickiewicz University, Institute of Physics Umułtowska 85, 61-614 Poznań, Poland

An effective Hamiltonian representing propagation and interaction of spin polarons forming in weakly doped antiferromgnets is derived. The systems becomes superconducting upon doping. Paring may be attributed to kinetic energy saving due to a particular type of motion of a spin bipolaron, which corresponds to creep of a line consisting of defects in the antiferromagnetic spin background which connects two holes. That line moves by shrinking and expanding at opposite ends. The kinetic energy term in the Hamiltonian is affective at all stages of this process, and the term related to the magnetic exchange does not have to intervene, not like in the case of a single propagating hole, which lowers the kinetic contribution to the energy. The effective attraction is strongest in the undoped system where the antiferromagnetic order is most robust, but the superconducting order parameter vanishes when the doping parameter decreases which should be attributed to emptying the spin polaron band and approaching the Mott insulator phase. Since the normal phase representing a gas of single spin polarons is unstable toward formation of bound hole pairs a pseudogap forms in the excitation spectrum.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : P. Wróbel

Address for correspondence : Institute for Low Temperature and Structure Research P. 0. Box 1410, 50-950 Wrocław 2, Poland

Email address : wrobel@int.pan.wroc.pl